### v1

# SaraCoder: Orchestrating Semantic and Structural Cues for Profit-Oriented Repository-Level Code Completion

Abstract
Retrieval-augmented generation (RAG) for repository-level code completion commonly relies on superficial text similarity, leading to results plagued by semantic misguidance, redundancy, and homogeneity, while also failing to resolve external symbol ambiguity. To address these challenges, we introduce SaraCoder, a Hierarchical Feature-Optimized retrieval framework. Its core Hierarchical Feature Optimization module systematically refines candidates by distilling deep semantic relationships, pruning exact duplicates, assessing structural similarity with a novel graph-based metric that weighs edits by their topological importance, and reranking results to maximize both relevance and diversity. Furthermore, an External-Aware Identifier Disambiguator module accurately resolves cross-file symbol ambiguity via dependency analysis. Extensive experiments on the challenging CrossCodeEval and RepoEval-Updated benchmarks demonstrate that SaraCoder significantly outperforms existing baselines across multiple programming languages and models. Our work proves that systematically refining retrieval results across multiple dimensions provides a new paradigm for building more accurate and robust repository-level code completion systems.
Figure 1: The pitfalls of pure similarity retrieval and the highlights of SaraCoder. Pink boxes illustrate traditional retrieval results based purely on surface similarity, while green boxes demonstrate results from our method SaraCoder.
Introduction
Code Large Language Models (CLLMs) (Izadi, Gismondi, and Gousios 2022; Li et al. 2022b; Allal et al. 2023) built on the Transformer architecture and trained on massive code datasets, have significantly empowered modern software development. They compress vast programming knowledge into billions or trillions of parameters, leading to successful deployments that streamline real-world development and boost productivity. However, as software projects grow in complexity, traditional Code LLMs struggle with repository-level completion. Their limited context windows cannot grasp long-range, cross-file dependencies, preventing a deep understanding of project semantics and architecture. This results in localized, often insufficient, suggestions. Directly processing entire repositories also causes inefficient retrieval, high latency, and an inability to leverage crucial higher-level software engineering knowledge.
To tackle these challenges, Retrieval-Augmented Generation (RAG) (Tang et al. 2023; Zan et al. 2022; Zhang et al. 2023) offers a compelling solution. RAG overcomes the limitations of traditional models by incorporating an external knowledge retrieval mechanism. Within this framework, an efficient retriever locates relevant code snippets, API documentation, or type definitions from the repository in real time. A generation model then integrates these retrieved results with the current context to create syntactically correct, semantically consistent, and standard-compliant code suggestions. RAG is progressively becoming a foundational technology for achieving accurate and trustworthy repository-level intelligent code completion.
Repository-level code completion can be divided into two key scenarios: In-File and Cross-File. In-File completion relies solely on the current file’s context. It is a fundamental and high-frequency task in repository-level intelligence. Cross-File completion involves dependencies on symbols from other files. The true complexity of modern software lies in these inter-module interactions. Cross-File completion is a crucial aspect of repository-level intelligence. However, some existing retrieval-augmented solutions rely too heavily on similarity search without proper filtering, and they often lack the crucial injection of symbolic relationships. This strategy exposes significant flaws in two types of scenarios, as shown in Figure 1: (1) In-File Scenarios: Similarity-based retrieval exhibits two critical flaws in practice. (a) Misleading Superficial Similarity occurs when retrievers recall code fragments that share surface-level characteristics but lack semantic relevance to the current context. The approach suffers from (b) Redundant Retrieval Homogenization, where duplicate or near-identical snippets are repeatedly retrieved, unnecessarily consuming valuable context window capacity that could be allocated for more diverse and informative content. (2) Cross-File Scenarios: Besides the limitations also observed in In-File scenarios, (c) External Symbol Ambiguity is a more severe limitation. Retrievers fail to capture critical dependencies such as unreferenced classes, interface constraints, or architectural rules, triggering error cascades including type inference failures, method signature mismatches, fabricated parameters/attributes, and architecture violations.
To address these challenges, we introduce SaraCoder, a Semantic-Aware Code Retrieval framework with Redundancy Reduction and Ambiguity Suppression for Repository-Level Code Completion. SaraCoder implements a hierarchical feature-optimized retrieval-augmented code completion framework, that goes beyond surface-level similarity by extracting deeper relationships such as control and data flow from code snippets. To further refine retrieval quality, we use semantic alignment distillation to capture deep semantic relationships. This is paired with a graph-based structural similarity metric that assesses a candidate’s structural proximity to the target context, weighting edit operations by their topological importance. To combat redundant retriveal homogenization, we integrate deduplication pruning and diversity-aware reranking, ensuring relevance while maximizing diversity. Furthermore, to resolve external symbol ambiguity of Cross-File scenarios in repository-level code completion, we add an external-aware identifier disambiguator that analyzes project-level dependencies. Our key contributions are:
  • We propose SaraCoder, a hierarchical and feature-optimized retrieval-augmented code completion framework. Designed for repository-level code completion, SaraCoder provides a comprehensive solution that prioritizes high-quality, diverse, and streamlined suggestions.
  • SaraCoder seamlessly integrates semantic alignment, redundancy pruning, topological proximity assessment, and diversity-driven reranking, resolving misleading superficial similarity and redundant retrieval homogenization. It also incorporates an external-aware identifier disambiguator that accurately resolves external symbol ambiguities by leveraging project-wide import dependencies.
  • SaraCoder’s coordination of semantic and structural signals enables strong performance even in large, resource-constrained codebases. Its design allows it to complement orthogonally other cross-file methods, delivering synergistic improvements when used in combination.
Related Work
Current RAG methods for repo-level code completion mainly rely on code similarity or cross-file dependencies.
Similar Code Snippet Retrieval for RAG
This approach enhances the quality of code LLM generation by retrieving semantically similar code snippets and integrating them into prompts, mimicking the reference behaviors of programmers. CodeSearchNet (Husain et al. 2020) pioneers large-scale code corpora construction, providing retrieval-based completion references; CodeRetriever (Li et al. 2022a) integrates pretrained models like CodeBERT (Feng et al. 2020) to enhance complex scenario handling; ReACC(Lu et al. 2022) combines vector and string-based retrieval to significantly optimize long-code processing; GraphCoder (Liu et al. 2024) improves code completion by using program dependencies for structured representations, allowing coarse-to-fine retrieval for Python and Java. However, like many other methods, its dependency analysis does not fully grasp deep semantic relationships in code. Additionally, most approaches rely too much on surface-level textual similarity. This often results in redundant retrieved content, wasting resources.
Cross-File Dependency Retrieval Augmentation
This method approaches code completion in complex repositories by leveraging cross-file code context (e.g., dependencies, dataflow, and subsequent similar code). Inspired by Ding et al.’s (Ding et al. 2023) observation that subsequent content of high-similarity snippets effectively informs completion, it injects these snippets into prompts. COCOMIC (Ding et al. 2024) dynamically fuses the context of this file with the cross-file entities retrieved by CCFinder (compressed into [SUM] vectors) through a joint attention mechanism to achieve location-aware code completion. DraCo (Cheng, Wu, and Hu 2024) extends this paradigm through dataflow-guided retrieval, parsing private repositories into code entities and constructing a repository-specific context graph reflecting dependencies. DraCo retrieves precise contextual knowledge from this graph to generate well-structured prompts, overcoming cross-file information barriers and repository-specific accuracy gaps. Current limitations include Python-exclusive implementation with type-sensitive dependencies lacking multilingual support.
Repository-Level Code Completion Evaluation
Although traditional code completion benchmarks (Chen et al. 2021; Austin et al. 2021) focus primarily on local file contexts and isolated code snippets, the inherent complexity of modern software development characterized by vast codebases and intricate in-file dependencies requires a more sophisticated evaluation approach. To address this gap, specialized benchmarks like RepoEval (Zhang et al. 2023), CrossCodeEval (Ding et al. 2023), RepoBench (Liu, Xu, and McAuley 2024), and ReccEval (Cheng, Wu, and Hu 2024) allow for evaluation of repository-level code completion. These establish standardized test scenarios spanning multiple languages, project scales, and contextual complexity levels to ensure rigorous and reproducible assessment. For example, RepoEval targets In-File completion; CrossCodeEval and ReccEval assess Cross-File completion requiring deep contextual understanding.
Figure 2: An illustration of SaraCoder framework. (1) Database Construction. This phase constructs a key-value codebase. This involves using a slicing algorithm to create induced graph slices, which are then precisely mapped to source code snippets. (2) Code Retrieval. This phase takes code context as input and retrieves similar code, then refines suggestions via Hierarchical Feature Optimization. Concurrently, an External-Aware Identifier Disambiguator clarifies external symbols via dependency analysis, delivering highly accurate candidates. (3) Code Generation. This phase generates prompts by integrating outputs from code retrieval with the code completion context. These prompts are then fed into an LLM to predict completion statements.
Method
As shown in Figure 2, SaraCoder is a hierarchical feature-optimized retrieval-enhanced code completion framework, applicable to both In-File and Cross-File completion scenarios. Formally, given a code context $C_{context}=\{x_{1},x_{2},\cdots,x_{n}\}$ and its containing file path $F$, the task aims to predict the next statement $\tilde{y}$.
Database and Initial Candidate Construction
To better represent code logic, we introduce a multi-level code context graph model that integrates control flow, data dependency, and control dependency (Liu et al. 2024). This structured representation offers enhanced generalization capabilities compared to serialization methods, enabling more effective capture of task-relevant context and facilitating easier adaptation to other languages. We utilize program slicing to generate precisely mapped, task-relevant subgraphs from source code on-demand, constructing structured codebases tailored to support specific analysis tasks. When a code completion task request occurs, we extract the Unfinished Code Context $C_{\text{context}}$ and the Import Statements $I$ from the code file . $C_{\text{context}}$ is then used to retrieve an initial candidate set of $top\_k\times p$  [^1] To ensure a high-quality final candidate set of top_ktop\_k results, we expand the initial candidate pool to $top\_k\times p$, allowing more candidates to participate in the Hierarchical Feature Optimization. code snippets $C$ via text similarity from structured codebases.
Hierarchical Feature Optimization (HF_OP)
Semantic Alignment Distillation (SAD)
SAD addresses Superficial Similarity Misguidance by leveraging the GraphcodeBERT (Guo et al. 2021), a pretrained model specialized in code understanding, to capture deep semantic relationships between code snippets. First, the query code $Q$ and candidate set $C$ are tokenized into subword sequences and uniformly padded or truncated to a fixed length $L=512$. Subsequently, during the feature encoding phase, a $768$-dimensional semantic vector $v_{s}$ is extracted for each code unit $s\in{Q}\cup C$, with vector space standardized through L2 normalization. When code repositories lack sufficient repetitive or relevant code, standard filtering methods are too strict, often leading to zero-candidate scenarios. This scarcity of reference material then hurts the accuracy of large language models. To fix this “one-size-fits-all” problem, we introduce a new dynamic quantile threshold mechanism. During the dynamic filtering phase, the cosine similarity set $S=\{\cos(v_{Q},v_{c})\mid c\in C\}$ is computed between the query vector $v_{Q}$ and all candidate vectors $v_{c}$. An adaptive threshold $\tau=\text{quantile}(S,0.75)$ is set at the 75th percentile, outputting filtered results $C_{SAD}=\{c\mid\cos(v_{Q},v_{c})\geq\tau\}$. To reduce redundant computation overhead and improve efficiency, a caching mechanism stores encoding results for high-frequency code.
Redundancy-Aware Pruning (RAP)
This module implements lightweight hash-based deduplication via exact text matching. Using the MD5 algorithm (Rivest 1992), it generates 128-bit hash fingerprints (single computation $\approx 0.02$ms, memory footprint $32$ bytes/hash) to eliminate verbatim duplicates from candidate set $C_{SAD}$ with minimal computational cost, significantly reducing downstream overhead. The module maintains a global hash set $H_{\text{seen}}$ to dynamically track processed sample fingerprints: for each candidate $c\in C_{SAD}$, if its MD5 hash $h_{c}\notin H_{\text{seen}}$, $c$ is added to deduplicated result set $C_{RAP}$ and $H_{\text{seen}}$ is updated. This achieves real-time processing with $O(N)$ time complexity. After SAD processing, the number of code snippets requiring MD5 hashing is limited and their structure is fixed by syntactic and semantic constraints. The MD5 collision resistance (theoretical probability $\approx 1.47\times 10^{-18}$) is sufficient for strict sensitivity. Additionally, MD5’s superior speed and lower memory footprint provide optimal cost-performance.
Topological Proximity Metric (TPM)
At this layer, the decaying subgraph edit distance (D-SED) is introduced to measure the graph similarity between the query graph $G_{q}$ and the candidate graph $G_{c}$ (Ranjan et al.; Zeng et al.).
$D-\operatorname{SED}\left(G_{q},G_{c}\right)=\sum_{op=\mathrm{O}}\gamma^{l(op)}\cdot c(op)$ (1)
Editing operations $O$ are the set of operations to transform $G_{c}$ to $G_{q}$, include adding, deleting, and modifying nodes and edges. Each operation $op\in O$ has a cost $c(op)$ and a hop count $l(op)$ from its “core node”. For simplicity, we choose the node with the largest ID as core node. Operations closer to the core exert greater structural influence. $\gamma\in(0,1)$ is an attenuation factor that reduces the cost weight for operations farther from the core node. After computing D-SED scores for each candidate $c\in C_{RAP}$, we compute a composite score $s$ as a weighted sum of text similarity (calculated during initial candidate generation) and structural similarity (D-SED scores). Subsequently, we generate $Q_{TPM}=[(\text{c},\text{s}),\ldots]$, ordered in descending score $s$.
Diversity-Aware Reranking (DAR)
This module implements a variability-aware ranking model based on the Maximal Marginal Relevance (MMR) (2017) algorithm to maximize result diversity while preserving relevance. It addresses homogeneity in traditional rankings through adversarial similarity calculation and dynamic weight adjustment.
[…]
$S$ contains items $(c,s)\in Q_{TPM}$ that have not been selected into $C_{final}$ yet. $\text{Sim}_{1}$ represents the relevance ($s_{i}=\pi_{2}\circ\iota_{c_{i}}(S)$) of item $c_{i}$ to query $q$. $\text{Sim}_{2}$ denotes the maximum cosine similarity between $c_{i}$ and any item $c_{j}$ in the selected set $C_{final}$. $\lambda$ is a trade-off parameter that balances the emphasis between relevance ($\text{Sim}_{1}$) and diversity ($\text{Sim}_{2}$).
Figure 3: Prompt template used in SaraCoder.
External-Aware Identifier Disambiguator (EAID)
This module enhances knowledge through external identifier augmentation. Firstly, file-level entity modeling parses code per file $F$, extracting method entities $E_{\text{method}}$ (functions/class methods with identifier, alias, line range $l_{\text{start}},l_{\text{end}}]$, parameter signature, scope) and class entities $E_{\text{class}}$ (class definitions with identifier, alias, line range $l_{\text{start}},l_{\text{end}}]$, member mappings) that built in the $F$. After that, it generates a structured identifier symbol table $ST_{\text{lib}}=\{\text{identifier}\mapsto\text{ syntax features}\}$, where $\mathit{identifier}$ corresponds to either: (1) the unique identifier of a method entity $\forall e\in E_{\text{method}}$, or (2) the unique identifier of a class entity $\forall c\in E_{\text{class}}$, with the mapped $\mathit{syntax\ features}$ containing all associated attributes for that entity. Subsequently, the dependency resolution mechanism processes all import statements ($I$) within the unfinshed file. For Intra-project Cross-module Reference, this phase retrieves complete entities ($E_{lib}$) from the pre-built entity library ($ST_{\text{lib}}$) by determining their corresponding file paths ($p$). These paths are constructed through decomposition of module components derived from either dotted names (e.g., my.module.MyClass) or relative imports (e.g., from .sub_module import MyClass), which are subsequently joined using directory separators (/) and appended with the .py file extension. For standard and third-party libraries, the system constructs a lightweight reference table $T_{\text{ext}}=\{\text{canonical name}\mapsto\text{alias}\}$ to efficiently manage external dependencies without full entity resolution. The enhanced prompts $P_{\text{E}}=I\oplus E_{lib}\oplus T_{\text{ext}}$.
Prompt Generation
Following code retrieval and external link resolution, SaraCoder employs an external LLM to generate subsequent statements. The final prompt $P_{\text{final}}$ is constructed by concatenating three components: the external symbols enhancement $P_{\text{E}}$ where entities are ordered by file import sequence reflecting call probability decay—with function entities populated with complete function bodies and class entities containing variable tables and method definitions; the similar code snippets $C_{final}$ containing code snippets strictly sorted in ascending order of similarity and annotated with source paths; the unfinished code context $C_{\text{context}}$. This architecture follows $P_{\text{final}}=C_{final}\oplus P_{\text{E}}\oplus C_{\text{context}}$. (Figure 3).
Experiments
Experimental Settings
To evaluate the performance of SaraCoder on repository-level code completion, we have formulated the following four research questions (RQs):
  1. RQ1
    Effectiveness in Cross-File Scenarios: How does SaraCoder perform when cross-context understanding is required, compared to other methods?
  2. RQ2
    Cost Analysis in Cross-File Scenarios: How does SaraCoder’s resource consumption compare to GraphCoder in Cross-File scenarios?
  3. RQ3
    Synergistic Gain Property: How does SaraCoder perform when integrated orthogonally with other cross-file methods?
  4. RQ4
    Advantage in In-File Scenarios: How does SaraCoder perform on tasks without cross-context requirements and what are its advantages?
Datasets
We primarily utilize two datasets here: CrossCodeEval and RepoEval-Updated. Table 1 shows details.
  • CrossCodeEval (Ding et al. 2023): This benchmark evaluates code completion in complex Cross-File scenarios like type inference and dependency analysis. It is ideal for assessing performance that requires a deep understanding of code across multiple files.
  • RepoEval-Updated (Liu et al. 2024): Expanded from RepoEval (2023), this new version, includes repositories of varying scales and duplication rates, offering a better way to evaluate In-File completion performance.
We use CrossCodeEval to test models on code completion tasks that involve complex cross-file dependencies. RepoEval-Updated will assess other aspects like basic syntax, common API usage, and local context understanding.
CrossCodeEval RepoEval-Updated
Python Java Python Java
Total Repositories 471 239 10 8
Total Files 1368 745 3258 8260
Total Task cases 2665 2139 2000 1600
Applicable Scenarios Cross-File completion In-File completion
Table 1: CrossCodeEval vs. RepoEval-Updated comparison.
Evaluation Indicators Setting
In this study, the following several evaluation indicators are used to assess the effect of code completion (Lu et al. 2021; Ding et al. 2023).
  • Code Exact Match (EM): Proportion of generated code exactly matching the ground truth. EM is given only for a perfect semantic and syntactic match.
  • Identifier Exact Match (ID_EM): The percentage of identifiers (variables, functions, etc.) perfectly matching the ground code. A high ID_EM means the model correctly understands the context and chooses the right names.
  • Identifier F1 Score (ID_F1): A more nuanced evaluation of identifier matching by combining precision and recall. This integration offers a more comprehensive assessment of identifier completion quality.
  • Edit Similarity (ES): Similarity metric between generated and ground-truth code based on edit distance. It tolerates slight variations, requiring the completed code to be highly similar in structure, syntax, and token order to the target.
Baseline Setting
We employ the following five methods as controls to evaluate the effectiveness of retrieval-augmented generation (RAG) in code completion: No RAG (Zero-shot-only baseline), Shifted RAG (Target-context dynamic retrieval), Vanilla RAG (Exemplar-similarity fixed retrieval), Repocoder (Iterative-fragment integration (Zhang et al. 2023)), Graphcoder (Structure-modeling CCG utilization (Liu et al. 2024)).
Language Methods Codegen2-7b Codegen25-7b deepseek-coder-6.7b-instruct CodeLlama-7b-Instruct
Code Match Identifier Match Code Match Identifier Match Code Match Identifier Match Code Match Identifier Match
EM ES EM F1 EM ES EM F1 EM ES EM F1 EM ES EM F1
Python No Rag 0.00 13.38 0.00 2.24 0.00 13.26 0.00 2.10 0.00 4.51 0.00 0.57 0.00 13.27 0.00 2.22
Shifted Rag 4.84 46.67 11.48 42.72 7.40 48.88 14.09 44.62 8.19 50.18 14.77 46.69 6.91 49.12 13.60 45.12
Vanilla Rag 9.48 50.97 17.15 47.81 12.39 53.92 23.61 53.14 13.00 54.00 26.63 55.77 11.45 52.93 19.23 50.49
Repocoder 12.47 54.08 21.57 51.89 16.62 56.80 25.73 57.85 17.11 58.11 26.71 56.46 15.14 56.28 24.56 54.22
Graphcoder 10.88 52.36 19.68 49.73 14.54 55.29 23.38 52.61 15.53 57.05 24.29 55.01 13.30 55.41 22.63 52.91
SaraCoder 15.04 56.03 24.44 54.68 18.36 58.30 27.28 56.22 19.72 59.93 28.52 58.26 17.91 58.37 27.77 56.82
Java No Rag 1.03 21.79 0.64 16.86 1.50 21.77 24.78 24.78 6.40 35.79 10.42 32.90 0.93 20.83 1.96 16.17
Shifted Rag 6.08 46.09 12.11 43.76 5.89 38.00 10.23 36.44 5.84 36.19 11.64 35.23 6.73 43.75 12.71 41.68
Vanilla Rag 9.30 47.42 15.71 45.69 10.38 40.76 15.29 39.91 8.88 33.59 15.01 33.51 10.93 45.01 17.81 44.08
Repocoder 10.71 41.83 16.18 41.51 12.16 42.38 17.63 41.69 9.58 34.25 15.76 34.13 13.23 46.01 19.87 45.14
Graphcoder 8.13 45.18 14.35 43.32 8.42 36.77 12.85 35.84 7.39 32.34 12.76 32.05 8.51 40.57 14.96 39.57
SaraCoder 11.73 46.69 18.37 45.47 11.40 39.33 16.22 38.97 11.92 34.55 17.72 34.99 12.95 42.71 19.54 42.33
Table 2: Performance comparison on the CrossCodeEval dataset. Numbers are shown in percentage ($\%$). The top results are bolded, and the second best are underlined.
Figure 4: Impact of top_k on CrossCodeEval. (The two on the left are Java tasks, and the two on the right are Python tasks.)
Model Selection
In this experiment, we select Codegen2-7b, Codegen25-7b (Nijkamp et al. 2023b, a), CodeLlama-7b-Instruct (Rozière et al. 2024) and deepseek-coder-6.7b-instruct (Guo et al. 2024) for code completion task inference.
Language Methods Codegen2-7b Codegen25-7b CodeLlama
Code Match Identifier Match Code Match Identifier Match Code Match Identifier Match
EM ES EM F1 EM ES EM F1 EM ES EM F1
Python No Rag 5.44 57.85 11.71 42.22 7.77 60.52 14.45 45.40 9.49 61.97 16.44 47.36
Shift Rag 4.87 58.36 11.64 42.91 7.44 60.20 14.17 44.78 6.95 60.35 13.75 45.36
Vanilla Rag 9.52 61.87 17.42 48.01 12.43 63.81 20.74 51.00 11.48 63.66 19.42 50.72
SaraCoder 12.16 54.16 18.37 45.47 11.50 44.90 16.22 38.95 13.32 48.04 19.54 42.34
Repocoder 12.50 64.48 21.87 52.09 16.66 66.67 25.99 55.06 15.19 66.24 24.74 54.39
Repocoder + SaraCoder 15.94 +3.44 66.43 +1.95 25.73 +3.86 54.80 +2.71 19.49 +2.83 68.53 +1.86 28.90 +2.91 57.49 +2.43 19.19 +4.00 68.45 +2.21 29.05 +4.31 57.47 +3.08
Draco 20.06 66.33 29.13 56.53 22.93 68.70 32.45 59.34 23.50 68.56 32.57 59.49
Draco + SaraCoder 24.06 +4.00 69.40 +3.07 34.00 +4.87 61.12 +4.59 27.05 +4.12 71.86 +3.16 36.80 +4.35 63.48 +4.14 27.20 +3.7 72.01 +3.45 37.29 +1.72 64.35 +4.85
Java No Rag 0.00 25.92 0.05 17.48 0.00 25.46 0.05 17.61 0.00 25.17 0.00 17.23
Shift Rag 6.45 54.84 12.11 43.75 6.08 44.73 10.27 36.46 7.11 50.96 12.72 41.68
Vanilla Rag 9.68 55.71 15.71 45.71 10.47 47.09 15.29 39.93 11.31 51.48 17.81 44.09
SaraCoder 12.16 54.16 18.37 45.47 11.50 44.90 16.22 38.95 13.32 48.04 19.54 42.34
Repocoder 11.22 56.89 17.72 47.41 10.85 47.93 16.18 41.50 13.60 52.17 19.87 45.14
Repocoder + SaraCoder 11.50 +0.28 56.09 -0.80 17.72 0.00 46.96 -0.45 11.27 +0.42 46.53 -1.40 16.41 +0.23 40.47 -1.03 14.26 +0.66 50.79 -1.38 20.62 +0.75 44.38 -0.76
Table 3: Performance benefits of SaraCoder when integrated orthogonally with other cross-file approaches ($\%$).
Language Methods Codegen2-7b Codegen25-7b deepseek-coder-6.7b-instruct CodeLlama-7b-Instruct
Code Match Identifier Match Code Match Identifier Match Code Match Identifier Match Code Match Identifier Match
EM ES EM F1 EM ES EM F1 EM ES EM F1 EM ES EM F1
Python No Rag 17.40 32.54 23.75 30.21 19.55 34.48 25.75 32.16 11.50 30.33 15.30 22.39 17.35 33.05 23.55 30.53
Shifted Rag 32.70 59.22 40.10 55.66 36.45 61.96 43.20 58.31 20.90 42.95 26.50 38.88 33.90 60.28 41.50 56.39
Vanilla Rag 38.70 63.58 46.45 60.43 42.25 66.26 48.75 62.79 22.20 41.48 27.85 37.58 40.30 65.03 47.55 61.06
Repocoder 37.60 61.98 45.10 58.47 40.55 64.48 46.85 60.71 21.35 40.18 26.65 35.93 39.60 63.71 47.05 59.78
Graphcoder 42.40 65.73 49.45 62.07 44.65 67.59 51.00 63.82 28.50 44.63 33.35 42.63 43.90 67.26 51.15 63.51
SaraCoder 42.60 65.92 50.15 62.61 44.50 67.79 51.10 63.84 28.25 46.91 33.45 42.95 45.00 68.27 52.00 63.97
Java No Rag 6.55 16.84 9.15 8.84 5.35 16.21 9.05 8.65 6.40 20.61 7.75 8.73 6.85 17.11 9.45 9.06
Shifted Rag 30.87 62.52 43.94 61.01 26.63 58.46 37.75 56.57 28.00 55.12 36.81 53.39 35.38 64.63 45.56 62.98
Vanilla Rag 33.50 63.82 45.44 62.08 32.00 61.52 41.56 59.48 21.13 46.51 31.19 45.15 38.56 66.46 48.06 64.90
Repocoder 30.13 60.01 42.31 57.10 28.75 57.73 37.88 54.46 22.19 46.93 32.06 44.26 35.38 62.33 44.44 59.63
Graphcoder 37.75 66.19 50.68 64.77 36.63 64.74 46.13 62.62 28.81 55.75 40.06 53.61 43.00 69.68 52.69 67.85
SaraCoder 37.93 67.06 50.93 65.48 36.75 65.54 46.44 62.61 29.75 56.38 40.75 54.18 42.88 69.37 52.00 67.37
Table 4: Performance comparison on the RepoEval-Updated dataset. Numbers are shown in percentage ($\%$). The top results are bolded, and the second best are underlined.
Main Results
For RQ1: Dominant Cross-File Code Accuracy.
Table 2 illustrates that SaraCoder surpasses the top-performing Repocoder on the CrossCodeEval dataset, achieving an average improvement of $1.50$ in EM, $0.77$ in ES, $1.11$ in ID_EM, and $0.61$ in ID_F1. This indicates SaraCoder provides more effective information and generates code with higher semantic accuracy, better capturing intended functionality. The enhanced ID_EM further shows SaraCoder’s superior ability to interpret context and select appropriate identifiers. These advancements effectively mitigate misleading superficial similarity and external symbol ambiguity, leading to more reliable and contextually relevant code. For Java code completion, SaraCoder shows better EM and ID_EM, with slightly lower ES and F1 scores. Because of Java’s static typing and verbosity, even small structural changes (e.g., bracket placement) can heavily affect ES/F1 scores, despite functional correctness.
For RQ2: Cost-Optimized Accuracy Advantage In Cross-File.
We experiment with code completion efficiency using codegen25-7b and Graphcoder. Our goal is to see how retrieving more similar cases (top_ktop\_k) impacts accuracy. Since using fewer top_ktop\_k cases saves input tokens [^2] For relevant explanations, please refer to the Appendix D.2, this study shows the balance between resources and accuracy. Our experiments demonstrate significant performance saturation for both retrieval methods when top_ktop\_k reaches 3-4, with no observable fluctuations upon increasing to top_k=10top\_k=10. SaraCoder achieves comprehensive superiority in Python tasks (e.g., 9.4% EM improvement) while maintaining advantages in Java tasks despite a marginal $0.1$ decrease in ES. Crucially, under resource-constrained top_k=1top\_k=1 conditions: all Python metrics outperform the baseline; three Java metrics (EM/ES/ID_EM) show improvements; and Java ID_F1 initially trails ($35.22$ vs. $35.27$) but ultimately surpasses the baseline at saturation ($35.84$ vs. $35.77$). Our method achieves performance breakthroughs at lower computational cost (stable at $top\_k\approx 4$) by reducing redundant and homogeneous cases (Figure 4).
For RQ3: Synergistic Integration of SaraCoder Achieves Enhanced Completion.
We examine two prominent methods that demonstrate exceptional performance in Cross-File scenarios. (1) Repocoder (Zhang et al. 2023), distinct from the original, assumes that if code snippets are similar, their subsequent content is also likely relevant. In the next search round, it specifically gets the code following those similar snippets (hereafter referred to as Repocoder). (2) Draco (Cheng, Wu, and Hu 2024), analyzes code to create entity dependency graphs, allowing detailed background knowledge retrieval. It then uses this information to create structured prompts. Currently, Draco only works with Python. As shown in Table 3, adding our method significantly boosts all four Python metrics (by $3.42$ to $4.52$) compared to using Repocoder or Draco alone. For Java, our method improves EM by 0.45 and ID_EM by $0.33$ over Repocoder, showing SaraCoder exhibits significant synergistic gain property with existing cross-file methods.  [^3] You can find the causes of synergistic gains in Appendix D.3.
Figure 5: Ablation study. (Each three-data-point group represents CodeGen2-7B, CodeGen2.5-7B, and CodeLlama-7B-Instruct models. Bar lengths show their average performance, with I-shaped error bars indicating S.D.)
For RQ4: Enhanced In-File Accuracy and Resource Efficiency.
On the RepoEval-Updated dataset (Table 4), SaraCoder shows superior semantic and identifier accuracy (surpassing the top-performing Graphcoder: +0.547+0.547 EM, +0.737+0.737 ES, +0.125+0.125 ID_EM, and +0.667+0.667 F1) for both Python and Java code completion. The cost analysis (Appendix D.1) further indicates SaraCoder generally performs better and exhibits higher stability across most Python metrics (excluding EM) and all Java metrics. This makes it particularly effective for resource-constrained environments, especially at lower top_ktop\_k values. However, SaraCoder’s gains over Graphcoder in code and identifier matching are smaller here than on CrossCodeEval. This is primarily because RepoEval-Updated projects contain a higher prevalence of similar code snippets, resulting in reduced code diversity. Overall, the conclusions align with those from the CrossCodeEval dataset.
Ablation Study
To understand the importance of each part of SaraCoder, we conduct ablation tests on the CrossCodeEval dataset (Figure 5). “-EAID” indicates disabling External-Aware Identifier Disambiguator, resulting in the loss of external dependency integration capabilities; “-HF_OP” denotes removing Hierarchical Feature Optimization, canceling the similar fragment screening mechanism; “-CCG” indicates disabling the code context graph, so it lost the understanding of code structure. The ablation experiments demonstrate that the complete SaraCoder achieves optimal performance, with all components positively contributing to repository-level completion. Notably, even without EAID, SaraCoder still outperforms Shift RAG and Vanilla RAG, and even surpasses Repocoder in Python tasks [^4] Detailed data can be found in Appendix, Table 4., proving that HF_OP screening substantially enhances case quality.
Conclusion and Outlook
This paper proposes SaraCoder, a profit-oriented, repository-level code completion method. It integrates semantic topology with disambiguation to tackle superficial similarity distraction, retrieval redundancy and rigidity, and external symbol ambiguity. The results of experiments on CrossCodeEval and RepoEval-Updated datasets show SaraCoder’s excellent accuracy and efficiency. However, though SaraCoder balances accuracy with reduced resource use by pruning unnecessary code examples, more optimal methods need further investigation.